OpenStudy (anonymous):

Suppose a pair of dice are rolled. Consider the sum of the numbers on the top of the dice and find the probabilities. (Enter the probabilities as fractions.) (a) 5, given that exactly one die came up 2.

OpenStudy (anonymous):

P(sum = 5) = favorable condition ( (3,2) or (2,3))/ total no. events(sample space) P(s=5) = 2/36 condition was given that one dice came up with 2 so only two cases will be there case 1 - dice1 = 2, dice2 = 3 case 2- dice1 = 3, dice2 = 2

OpenStudy (anonymous):

I tried that, but the answer wasn't right :/

OpenStudy (anonymous):

P(s=5) = 2/36 = 1/18